A bleaching solution contains sodium hypochlorite, NaClO, dissolved in water. The solution is 0.650 m NaClO. What is the mole fraction of sodium hypochlorite?


1
Expert's answer
2020-11-02T09:02:51-0500

n(H2O)=m/M=1000/18=55.556(mol)

n(NaClO)=0.650(mol)

ntotal=55.556+0.650=56.206(mol)

X=n(NaClO)/ntotal=0.650/56.206=0.012